# Price formation without fuel costs: the interaction of elastic demand with storage bidding

### v1

@@ 3.1 Implementation details
To demonstrate how price formation works, modeling is done for three European countries (Germany, Spain, and the United Kingdom) using 70 years of historical hourly weather data from 1951 to 2020 [48]. We use 70 years of data for different countries to capture a wide variety of weather conditions, rather than illustrating a 70-year pathway of system transformation. Systems are modeled with only a limited selection of technologies: onshore wind, utility-scale solar photovoltaics, batteries for short-term storage, and a hydrogen-based chemical storage medium for long-term storage (e.g. underground hydrogen storage). The techno-economic parameters, such as investment costs and efficiencies, are provided in Table S1. The figures in the main body show results for Germany, while the results for Spain and the United Kingdom are provided in D.
The modeling framework used is Python for Power System Analysis (PyPSA) [49], and the open source code to reproduce our experiments is available on GitHub; [^4] https://github.com/fneum/price-formation (v0.1.0) output data was deposited on Zenodo. [^5] https://doi.org/10.5281/zenodo.12759248 (v0.1.0) The experiments were run on a high-performance computing cluster with PyPSA v0.27.1 [49], linopy v0.3.8 [50], and Gurobi v11.0.2 [45]. For the resulting linear (LP) and quadratic (QP) problems, we used the interior point method without crossover, a barrier convergence tolerance of $10^{-7}$, and 32 threads per model. The maximum memory requirement per scenario was 22 GB.
In this setup, we then vary the number of weather years in the simulation, the choice of the demand curve (Section 3.2), long-term optimization with capacity expansion versus short-term optimization with fixed capacities, capacity perturbations of ±$\pm 5\%$ around the long-term equilibrium, and myopic versus perfect foresight dispatch (Section 3.3).
3.2 Demand modeling
@@
The demand curve is varied between a perfectly inelastic demand, an inelastic step function up to VOLL, a linear demand curve, and a piecewise-linear (PWL) approximation of a log-linear demand curve, as shown in Fig. 2. Non-linear curves are not considered because solving becomes inefficient once we go beyond quadratic terms in the objective function. The same demand curve is used for each hour to isolate effects of supply from demand. The level of demand is chosen to represent a token average of 100 MW in each country to simplify case comparisons.
The simplest case is demand that is completely inelastic up to a given value of lost load (VOLL). For a given demand $0\leq d_{c}\leq D_{c}$ with value of lost load $V_{c}$, the utility function is
$U_{c}(d_{c})=V_{c}d_{c}$ (8)
The demand curve is a step function. It is perfectly inelastic up to a price of $p_{c}=U_{c}^{\prime}(d_{c})=V_{c}$, at which point it is perfectly elastic. Implementing this demand function in the optimisation problem Eqs. 1 and 2 results in a linear problem (LP). We model a single consumer $c$ with $V_{c}=2000$/MWh and peak load $D_{c}=100$ MW.
@@
$p_{c}=U_{c}^{\prime}(d_{c})=a_{c}-b_{c}d_{c}$ (10)
This results in a quadratic problem (QP), similar to one of the approaches in De Jonghe et al. [42]. For a linear demand curve, the elasticity varies along it, becoming more elastic at higher prices.
To approximate a demand curve of log-linear form $\ln(p_{c})=a_{c}-b_{c}\ln(d_{c})$ with constant price elasticity, we use an aggregated demand curve given by several demands $d_{c}\in\left[0,D_{c}\right]$ with different constants ($a_{c},b_{c}$), so that we get a piecewise-linear curve with different slopes. In Fig. 2, an example approximation is shown with three linear segments. Here, we use three segments with parameters $a=(8000,400,200)^{\top}$, $b=(80,40,20)^{\top}$, and $D=(95,5,10)^{\top}$. For sensitivity cases with higher and lower elasticity, coefficients $a$ and $b$ are halved or doubled, respectively.
@@ 3.3 Myopic dispatch
Assuming perfect foresight, where future periods of scarcity and abundance are fully anticipated, is unrealistic. The main challenge for running the system more realistically with myopic foresight, e.g., only 2-4 days, is dispatching the long-duration storage so that it charges in periods of excess for periods of scarcity outside the foresight horizon.
To evaluate the impact of myopic foresight on price formation for different demand modeling cases, we initially derive optimized capacities from an LT model trained on 35 randomly selected years from the 70-year dataset. Subsequently, we run the dispatch in an ST model using the remaining 35 years, comparing cases with perfect foresight with those with myopic foresight. We also apply perturbations of ±$\pm 5\%$ to all generation, conversion and storage capacities.
For cases with myopic foresight, we apply a rolling-horizon dispatch optimization with a 96-hour foresight and a 48-hour overlap between windows. This approach aligns with the horizon of current forecast skill and provides sufficient information for the short-term battery operation [51].
@@
A similar methodology was employed in [35], utilizing a state-of-charge-dependent storage value, as also noted in [21]. In contrast, we maintain a simpler method with a constant MSV to demonstrate that reasonable results can be achieved even with this simplification. This assumption could be justified if industrial offtake users with long-term contracts and producers on global markets contribute to stabilizing the hydrogen price.
@@ 4 Results
Figure 3: Different price duration curves: Panels a) and b) show electricity price duration curves as the number of years used in the optimization rises from 1 to 70. Panels c) and d) show the corresponding hydrogen price duration curves for the same simulations. Panels e) and f) show for the 70-year optimization how the prices change when going from the long-term (LT) to short-term (ST) model with the LT capacities, and then perturbing the asset capacities in the ST model up and down by 5%.
@@ 4.1 Results from the long-term model
Fig. 3 shows different price duration curves from the model with inelastic demand up to a VOLL of 2000 €/MWh (left column) and elastic demand with a piecewise-linear approximation of a -5% price elasticity of demand (right column). The results assume hydrogen is used for long duration storage.
Panels a) and b) in Fig. 3 reveal how the electricity price duration curves from the long-term model vary as the number of years in the simulation is increased from one year up to the full 70 years. For a duration curve to be meaningful, it should not change significantly as the number of years increases, but for the inelastic case, it becomes more and more singular as more years are considered. This is because high prices are clustered into the wind-scarce years, which drive the investment (Fig. S9). For the elastic case, the elastic demand curve spreads the scarcity pricing out over the years, resulting in price duration curves that have a more consistent shape as the number of years in the model is varied Figs. 4 and 5). Furthermore, going from inelastic to elastic demand for 70 years of data, the fraction of zero-price hours reduces from 89% to 31%, and the fraction of hours with prices over 400 €/MWh reduces from 8.5% to 3.9% so that the extreme variability of prices is substantially damped. This means that revenues for most technologies are more evenly distributed across different price bands (Fig. S7) instead of being concentrated in just a few high-priced hours. These results reproduce observations from sector-coupled models [17, 20], where elastic demand plays the same role as demand flexibility to reduce and stabilize prices.
We conclude from this that price duration curves from models with inelastic demand are not robust enough between simulations that anything can be deduced from them. Strong price bifurcations seen in previous studies (e.g. [3]) are modelling artefacts that cannot be taken at face value, since they are not consistent as the simulation length is changed and they depend strongly on how demand is modelled.
Panels c) and d) in Fig. 3 show hydrogen’s corresponding price duration curves. Since hydrogen can be inexpensively stored in underground caverns ($\approx$ 0.15 €/kWh, cf. Table S1), its price is more stable between different hours. This is because between periods of scarcity there is no system benefit in moving hydrogen between hours without binding capacity constraints. However, it still shows variations between wind-scarce and windy years (Fig. S9). The price variability is much more extreme for the inelastic case, mirroring the tight coupling between electricity prices and marginal storage values.
@@ 4.2 Results from short-term model with perfect foresight
Panels e) and f) in Fig. 3 show the results of taking the optimal capacities from the long-term (LT) model, fixing them, and rerunning the operational short-term (ST) optimization of the model with the same weather years and perfect operational foresight. We then perturb all capacities of generation and storage up and down by 5% to reveal how sensitive the prices are to capacities outside the long-term equilibrium.
Going from LT to ST model, we see the result of the theoretical argument in Section 2.3 confirmed: the prices are identical for an elastic demand, where we have a unique mapping from demand to price, but for the inelastic case the lack of price regularization causes the prices of the ST model to diverge from the LT prices. Furthermore, prices are susceptible to small capacity perturbations, either becoming very high in many hours if capacity is tight or collapsing to zero if capacities are slightly over-dimensioned. In the elastic case, the price changes follow the same direction but are milder in magnitude and more regular over time. Average prices rise by 104% with 5% less capacity and go down by 63% with 5% more capacity (Table S2), incentivizing market exits or entries with investments towards the long-term equilibrium.
If some dispatchable capacity is retained (Fig. S2) or reserve power capacity for re-electrification is procured outside the energy-only market (Fig. S3), our results remain largely unchanged. Some high-price episodes gradually disappear, both for inelastic and elastic demand, as scarcity costs of power deficits are replaced by the cost of reserve capacities or dispatchable power plants. For the elastic case, we still see identical prices between the LT and ST models.
@@ 4.3 Inter-annual variability
To explore this, we show in Figs. 4 and 5 the distribution of average annual and monthly baseload prices from the optimized 70-year period. The inelastic demand cases show a very high variance with a strong clustering of high prices in the 1960s. In contrast, prices cluster ever tighter for increasing elasticity and are more evenly distributed. In all cases except for fully inelastic demand, we see a seasonal pattern with higher prices in winter and broadly lower prices in summer. If we compare the results to the distribution of prices over the last 19 years [53, 54], we see a comparable range as seen in the conventional system due to fossil fuel price variability. [^8] It should be noted that the historical wholesale market prices do not yield full cost recovery for the assets in the system because aspects like renewable subsidies are not included. Such stable prices would increase certainty and reduce risk for investors.
@@ 4.4 Results from short-term model with myopic foresight
Figure 6: Price duration curves comparing short-term models with myopic foresight on unknown years, i.e. the LT and ST model take two disjunct sets of 35 weather years as input data. Blue is from the original LT model based on 35 years of data. Orange is from a ST model with optimized capacities from LT model and perfect foresight on 35 unseen test years. Green is another ST model with myopic foresight of 96 hours and a rolling horizon overlap of 48 hours for the same unseen test years. Dotted lines add capacity perturbations of ±$\pm 5\%$ to that.
@@
Fig. 6 compares price duration curves from myopic model runs (96-hour foresight, 48-hour overlap) with the LT and ST models with perfect foresight. Since the MSV is set to a constant value, there are now only two large non-zero steps in the price duration curves corresponding to the charging and discharging bids for hydrogen storage. The small intermediary steps are caused by interactions with the bids of the battery storage and the elastic demand, if present (Fig. S1).
In the inelastic case, there is a large increase in hours with load-shedding going from the LT capacity expansion model to the ST operational model, since the weather years of the ST model were different to those in the LT model. With myopic foresight the heuristic storage bidding based on a constant MSV acts to regularise prices, resulting in two main non-extreme price levels besides load-shedding and zero-price hours. A constant MSV was not something found in the LT model (see Fig. 3, panel c)) but exogenously set in the model inspired by observations from the elastic case (see Fig. 3, panel d)). Moreover, the monthly average baseload price distribution is similarly even as with demand elasticity (Fig. S4). Taking the constant averaged MSV blends out the volatility of the MSV formation in the absence of demand elasticity. With myopic foresight, the peak load shedding increases by 13.9% from 77.6 MW to 88.3 MW, while the average load served is just reduced by 0.5% from 99.96 MW to 99.47 MW. To compare the system operation between perfect and myopic foresight with and without elastic demand, we refer to Fig. S8.
In the elastic case, the price duration curves are similar. Unlike in Fig. 3, the price duration curves between the LT and ST models are close but not identical. This is again because the wind and solar capacity factors are taken from different weather years. The minor differences observed between the ST models with myopic and perfect foresight outline where the heuristic storage bids under or overestimate the MSV. With myopic foresight, the average electricity price increases from 95.4 €/MWh to 131.7 €/MWh. This reflects the cost of lacking foresight. Peak demand curtailment increases by 23.2%, whereas the average load served is reduced by only 0.8%.
@@
These results highlight that long foresight horizons are not required for operating a highly renewable energy system with storage. A horizon of a few days corresponding to a current forecasting skill suffices for battery scheduling and simple bidding heuristics for long-duration energy storage yield acceptable dispatch decisions.

### v2

@@ 3.1 Implementation details
To demonstrate how price formation works, modelling is done for three European countries (Germany, Spain, and the United Kingdom) using 70 years of historical hourly weather data from 1951 to 2020 [53]. We use 70 years of data for different countries to capture a wide variety of weather conditions, rather than illustrating a 70-year pathway of system transformation. Systems are modeled with only a limited selection of technologies: onshore wind, utility-scale solar photovoltaics, batteries for short-term storage, and a hydrogen-based chemical storage medium for long-term storage (e.g. underground hydrogen storage). The techno-economic parameters, such as investment costs and efficiencies, are provided in Table S1. The figures in the main body show results for Germany, while the results for Spain and the United Kingdom are provided in F.
The modelling framework used is Python for Power System Analysis (PyPSA) [54], and the open source code to reproduce our experiments is available on GitHub; [^5] https://github.com/fneum/price-formation (v0.2.0) output data were deposited on Zenodo. [^6] https://doi.org/10.5281/zenodo.12759247 (v0.2.0) The experiments were run on a high-performance computing cluster with PyPSA v0.27.1 [54], linopy v0.3.8 [55], and Gurobi v11.0.2 [49]. For the resulting linear (LP) and quadratic (QP) problems, we used the interior point method without crossover, a barrier convergence tolerance of $10^{-7}$, and 32 threads per model. The maximum memory requirement per scenario was 22 GB.
In this setup, we then vary the number of weather years in the simulation, the choice of the demand curve (Section 3.2), long-term optimisation with capacity expansion versus short-term optimisation with fixed capacities, capacity perturbations of ±$\pm 5\%$ around the long-term equilibrium, and myopic versus perfect foresight dispatch (Section 3.3).
3.2 Demand modelling
@@
A single demand curve is modelled to represent all aggregated consumers. The demand curve is varied between a perfectly inelastic demand, an inelastic step function up to VOLL, a linear demand curve, and a piecewise-linear (PWL) approximation of a log-log demand curve, as shown in Fig. 2. For most of the modelling analysis we focus on the case of inelastic demand up to VOLL, since this is the common modelling approach and also reflects the price caps enforced in many markets, and the PWL log-log elasticity case, since this was empirically measured in the German market [48]. Non-linear curves are not considered because solving becomes inefficient once we go beyond quadratic terms in the objective function. The same demand curve is used for each hour to isolate effects of supply from demand. The level of demand is chosen to represent a token average of 100 MW in each country to simplify case comparisons.
The simplest case of perfectly inelastic demand is modelled by removing the utility from the objective function, since it is effectively infinite, and replacing the demand variable by a demand constant.
The next simplest case is that of demand that is completely inelastic up to a given value of lost load (VOLL). For a given demand $0\leq d\leq D$ with value of lost load $V$, the utility function is
$U(d)=Vd$ (8)
The demand curve is a step function. It is perfectly inelastic up to a price of $p=U^{\prime}(d)=V$, at which point it is perfectly elastic. Implementing this demand function in the optimisation problem Eqs. 1 and 2 results in a linear problem (LP). For the modelling we assume a VOLL of $V=2000$/MWh and a peak load of $D=100$ MW.
@@
$p=U^{\prime}(d)=a-bd$ (10)
This results in a quadratic problem (QP), similar to one of the approaches in De Jonghe et al. [45]. For a linear demand curve, the elasticity varies along it, becoming more elastic at higher prices. Using parameters $a=2000$ and $b=20$ results in an elasticity of $-5\%$ near the approximate average system price $p=100$/MWh.
To approximate a demand curve of log-log form $\ln(p)=a-b\ln(d)$ with constant price elasticity, we use an aggregated demand curve given by several demands $d\in\left[0,D\right]$ with different constants ($a,b$), so that we get a piecewise-linear curve with different slopes. In Fig. 2, an example approximation is shown with three linear segments. Here, we use three segments with parameters $a=(8000,400,200)^{\top}$, $b=(80,40,20)^{\top}$, and $D=(95,5,10)^{\top}$, which results in an elasticity of -5% at a price of 100 €/MWh and a demand around 100 MW. The elasticity of $-5\%$ reflects the empirical analysis of the German market made in [48]. For sensitivity cases with higher and lower elasticity at the same price levels, coefficients $a$ and $b$ are halved or doubled which results in elasticities of $-10\%$ and $-2.5\%$ respectively at a price of 100 €/MWh.
In our main scenarios, we assume for simplicity that there are no inter-temporal dependencies for the demand elasticity, i.e. the demand level $d_{t}$ at time $t$ only depends on the price $p_{t}$. This is consistent with how the empirical measurements of German elasticity were made, since the inter-temporal effects were already present in the data [48]. In reality there may be complex dependencies on the hours before or after. In particular, some loads may not be able to reduce their demand for multiple consecutive hours, or they may have to increase the demand later to recover the lost demand. This can be represented in the model with cross-price elasticities between different hours [45], which adds bilinear terms between the hourly demands in the objective and does not affect our theoretical analysis, or by representing load shifts using a virtual storage for the demand, which is a common approach for e.g. flexible electric vehicle charging or heat pump operation [56]. In B, we show in a sensitivity analysis that cross-elastic terms have only a limited effect on our results.
@@ 3.3 Myopic dispatch
Assuming perfect foresight over multiple years, where future periods of scarcity and abundance are fully anticipated, is unrealistic. While forecasts may be reliable 1-3 days ahead, and hydroelectric planners are accustomed to making seasonal forecasts for water levels in their dams, the multi-week time scales on which wind power vary are harder to anticipate in advance [57]. In particular this creates challenges for the dispatch of long-duration storage so that it charges during days of excess in order to discharge during periods of scarcity outside the myopic foresight horizon.
To evaluate the impact of myopic foresight on price formation for different demand modelling cases, we initially derive optimised capacities from an LT model optimised using 35 randomly selected years from the 70-year dataset (D). Subsequently, we run the dispatch in an ST model using the remaining 35 years, comparing cases with perfect foresight with those with myopic foresight. We also apply perturbations of ±$\pm 5\%$ to all generation, conversion and storage capacities, to show the impacts of imperfectly anticipating the system’s needs.
For cases with myopic foresight, we apply a rolling-horizon dispatch optimisation with a 96-hour look-ahead horizon and a 48-hour overlap between windows. This approach aligns with the horizon of current forecast skill and provides sufficient information for the short-term battery operation [57]. Since there are already good weather forecasts up to 5 days ahead [57, 58], this represents a conservative worst case for the system dispatcher.
@@
A similar methodology was employed in [38], utilising a state-of-charge-dependent storage value, as also noted in [24]. In contrast, we maintain a simpler method with a constant MSV to demonstrate that reasonable results can be achieved even with this simplification. This assumption could be justified if industrial offtake users with long-term contracts and producers on global markets contribute to stabilising the hydrogen price.
@@ 4 Results
Figure 3: Different price duration curves: Panels a) and b) show electricity price duration curves as the number of years used in the optimisation rises from 1 to 70. Panels c) and d) show the corresponding hydrogen price duration curves for the same simulations. Panels e) and f) show for the 70-year optimisation how the prices change when going from the long-term (LT) to short-term (ST) model with the LT capacities, and then perturbing the asset capacities in the ST model up and down by 5%.
@@ 4.1 Results from the long-term model
Fig. 3 shows different price duration curves from the model with perfectly inelastic demand up to a value of lost load (VOLL) of 2000 €/MWh (left column) and with a piecewise-linear (PWL) approximation of -5% price elasticity of demand (right column). The results assume hydrogen is used for long duration storage.
Panels a) and b) in Fig. 3 reveal how the electricity price duration curves from the long-term model vary as the number of years in the simulation is increased from one year up to the full 70 years (D). For a duration curve to be meaningful, it should not change significantly as the number of years increases, but for the perfectly inelastic case, it becomes more and more singular as more years are considered. This is because high prices are clustered into the wind-scarce years, which drive the investment (Fig. S12). As the high prices cluster in the worst weather years, prices at other times collapse, since at these times there is over-capacity. For the case with demand elasticity, the elastic demand curve spreads the scarcity pricing out over the years, resulting in price duration curves that have a more consistent shape as the number of years in the model is varied. Furthermore, going from the perfectly inelastic to the price-responsive case for 70 years of data, the fraction of zero-price hours reduces from 89% to 31%, and the fraction of hours with prices over 400 €/MWh reduces from 8.5% to 3.9% so that the extreme variability of prices is substantially damped. This means that revenues for most technologies are more evenly distributed across different price bands (Fig. S9) instead of being concentrated in just a few high-priced hours. These results reproduce observations from sector-coupled models [20, 23], where demand elasticity plays the same role as demand flexibility to reduce and stabilise prices.
We conclude from this that price duration curves from models with perfectly inelastic demand are not robust enough between simulations that anything can be deduced from them. Strong price bifurcations seen in previous studies (e.g. [4]) are modelling artefacts that cannot be taken at face value, since they are not consistent as the simulation length is changed and they depend strongly on how demand is modelled.
Panels c) and d) in Fig. 3 show hydrogen’s corresponding price duration curves. Since hydrogen can be inexpensively stored in underground caverns ($\approx$ 0.15 €/kWh, cf. Table S1), its price is more stable between different hours. This is because between periods of scarcity there is no system benefit in moving hydrogen between hours without binding capacity constraints. However, it still shows variations between wind-scarce and windy years (Fig. S12). The price variability is much more extreme for the perfectly inelastic case, mirroring the tight coupling between electricity prices and marginal storage values.
@@ 4.2 Results from short-term model with perfect foresight
Panels e) and f) in Fig. 3 show the results of taking the optimal capacities from the long-term (LT) model, fixing them, and rerunning the operational short-term (ST) optimisation of the model with the same weather years and perfect operational foresight. We then perturb all capacities of generation and storage up and down by 5% to reveal how sensitive the prices are to capacities away from the long-term equilibrium.
Going from the LT to the ST model, we see the result of the theoretical argument in Section 2.3 confirmed: the prices are identical for demand elasticity, where we have a unique mapping from demand to price, but for the perfectly inelastic case the lack of price regularisation causes the prices of the ST model to diverge from the LT prices. Furthermore, prices are susceptible to small capacity perturbations, either becoming very high in many hours if capacity is tight or collapsing to zero if capacities are slightly over-dimensioned. In the elastic case, the price changes follow the same direction but are milder in magnitude and more regular over time. Average prices rise by 104% with 5% less capacity and go down by 63% with 5% more capacity (Table S4), incentivising market exits or entries with investments towards the long-term equilibrium.
4.3 Forcing dispatchable capacity to limit load reduction
Demand elasticity means that there are more regular reductions in demand, as can be seen from the load duration curves in Fig. S6. This can be a problem for some consumers if the reductions last more than a few hours. In our results there are reductions of 20% lasting 1-2 days roughly every other year during dark wind lulls, while there are 2 events in the 70 year sample lasting 10 days. To mitigate these high price events, we explore in Fig. S4 how forcing dispatchable capacity affects the model.
If some conventional dispatchable capacity is retained (Fig. S3) or power capacity for re-electrification of hydrogen is procured outside the energy-only market (Fig. S4), our results remain largely unchanged. Some high-price episodes gradually disappear, both for the inelastic and elastic cases, as scarcity costs of power deficits are replaced by the cost of reserve capacities or dispatchable power plants. For the elastic case, we still see identical prices between the LT and ST models.
@@ 4.4 Inter-annual variability
To explore this, we show in Figs. 4 and 5 the distribution of average annual and monthly baseload prices from the optimised 70-year period. The perfectly inelastic demand cases show a very high variance with a strong clustering of high prices in the 1960s. In contrast, prices cluster ever tighter for increasing elasticity and are more evenly distributed. In all cases except for fully inelastic demand, we see a seasonal pattern with higher prices in winter and broadly lower prices in summer. If we compare the results to the distribution of prices over the last 19 years [60, 61], we see a comparable range as seen in the conventional system due to fossil fuel price variability. [^9] It should be noted that the historical wholesale market prices do not yield full cost recovery for the assets in the system because aspects like renewable subsidies are not included. Such stable prices would increase certainty and reduce risk for investors, as annual revenues would also be less volatile (Fig. S10).
@@ 4.5 Results from short-term model with myopic foresight
Figure 6: Price duration curves comparing short-term models with myopic foresight on unknown years, i.e. the LT and ST model take two disjunct sets of 35 weather years as input data. Blue is from the original LT model based on 35 years of data. Orange is from a ST model with optimised capacities from the LT model and perfect foresight on 35 unseen test years. Green is another ST model with myopic foresight of 96 hours and a rolling horizon overlap of 48 hours for the same unseen test years. Dotted lines add capacity perturbations of ±$\pm 5\%$ to that.
@@
Fig. 6 compares price duration curves from myopic model runs (96-hour foresight, 48-hour overlap) with the LT and ST models with perfect foresight. Since the MSV is set to a constant value, there are now only two large non-zero steps in the price duration curves corresponding to the charging and discharging bids for hydrogen storage. The small intermediary steps are caused by interactions with the bids of the battery storage and the demand curve, if present (Fig. S2).
In the perfectly inelastic case, there is a large increase in hours with load-shedding going from the LT capacity expansion model to the ST operational model, since the weather years of the ST model were different to those in the LT model. With myopic foresight the heuristic storage bidding based on a constant MSV acts to regularise prices, resulting in two main non-extreme price levels besides load-shedding and zero-price hours. A constant MSV was not something found in the LT inelastic model (see Fig. 3, panel c)) but exogenously set in the model inspired by observations from the elastic case (see Fig. 3, panel d)). As a result, the monthly average baseload price distribution has a similar regularity to the case with demand elasticity (Fig. S5). Taking the constant averaged MSV blends out the volatility of the MSV formation in the absence of demand elasticity. With myopic foresight, the peak load shedding increases by 13.9% from 77.6 MW to 88.3 MW, while the average load served is just reduced by 0.5% from 99.96 MW to 99.47 MW. To compare the system operation between perfect and myopic foresight with and without demand elasticity, we refer to Fig. S11.
In the elastic case, the price duration curves are similar. Unlike in Fig. 3, the price duration curves between the LT and ST models are close but not identical. This is again because the wind and solar capacity factors are taken from different weather years. The minor differences observed between the ST models with myopic and perfect foresight outline where the heuristic storage bids under or overestimate the MSV. With myopic foresight, the average electricity price increases from 95.4 €/MWh to 126.2 €/MWh. This reflects the cost of lacking foresight. Peak demand curtailment increases by 23.2%, whereas the average load served is reduced by only 0.8%. The drop in total welfare is just 0.1%, since the welfare is dominated by the demand utility (see Table S5).
@@
These results highlight that long foresight horizons are not required for operating a highly renewable energy system with storage. A horizon of a few days corresponding to current forecasting skill suffices for battery scheduling and simple bidding heuristics for long-duration energy storage yield acceptable dispatch decisions. Any estimates of weather conditions further ahead will only help to improve the economic dispatch. A live website was developed by one of the authors to demonstrate how 24-hour-ahead forecasts combined with fixed hydrogen MSV are able to dispatch a fully renewable German power system over 10 years of weather data [62].